Projection Close 1/25
To do this evening/ was completed: In the morning I'll be monitoring the snow/metro and bus service.
Inspected all of Incredible Shrinking Man, which is in good condition despite crappy leader on most reels. Photos up on the drive.
Presentation Issues:
Preshow Clips/Playlist:
Trailer Updates:
Special Events/ AV set up:
Ordering/Receiving (posters & trailers):
Ordering/Receiving (content): Managed to get six more of the NoBudge shorts DCP'd this evening, so there's only one more to go! Please DCP Goodnight tomorrow, then we can ingest the rest. As Alec mentioned, the video files and DCPs are in the NoBudge 49 folder on the desktop.
Preshow Clips/Playlist:
Trailer Updates:
Special Events/ AV set up:
Ordering/Receiving (posters & trailers):
Ordering/Receiving (content): Managed to get six more of the NoBudge shorts DCP'd this evening, so there's only one more to go! Please DCP Goodnight tomorrow, then we can ingest the rest. As Alec mentioned, the video files and DCPs are in the NoBudge 49 folder on the desktop.
Keys for Possessor are in.
Outgoing/Shipping (content):
Print/video Inspection: Incredible Shrinking Man is on house reels now after inspection. I carried on with Jurassic Park, which can get finished up tomorrow.
Asset tracking doc updated/checked?:
Maintenance/cleaning: There is a gnarly smell in T7 and we considered switching No Other Choice (last round) to a different theatre, but we kept everything as is per Matthew and Jess' suggestion. Turns out, it is likley that the smell is a gas leak. We'll keep y'all updated.
Outgoing/Shipping (content):
Print/video Inspection: Incredible Shrinking Man is on house reels now after inspection. I carried on with Jurassic Park, which can get finished up tomorrow.
Asset tracking doc updated/checked?:
Maintenance/cleaning: There is a gnarly smell in T7 and we considered switching No Other Choice (last round) to a different theatre, but we kept everything as is per Matthew and Jess' suggestion. Turns out, it is likley that the smell is a gas leak. We'll keep y'all updated.
The lamp in T2 did turn off properly during shutdown this evening.
Cue sheets updated/checked (every shift): Yes.
Marquee updated?:
Posters updated?:
Cleaning lights on for the night?:
Projection cart by entrance?:
Additional Notes:
Cue sheets updated/checked (every shift): Yes.
Marquee updated?:
Posters updated?:
Cleaning lights on for the night?:
Projection cart by entrance?:
Additional Notes:
Comments
Post a Comment